# Towards Timetronics with Photonic Systems

### v1

@@ Abstract
Periodic driving of systems of particles can create crystalline structures in time. Such systems can be used to study solid-state physics phenomena in the time domain. In addition, it is possible to engineer the wave-number band structure of optical systems and to realize photonic time crystals by periodic temporal modulation of the material properties of the electromagnetic wave propagation medium. We introduce here a versatile averaged-permittivity approach which empowers emulating various condensed matter phases in the time dimension in a traveling wave resonator. This is achieved by utilizing temporal modulation of permittivity within a small segment of the resonator and the spatial shape of the segment. The required frequency and depth of the modulation are experimentally achievable, opening a pathway for research into the practical realisation of crystalline structures in time utilising microwave and optical systems.
Electromagnetic waves can propagate in media where the refractive index changes periodically in space or time [1, 2, 3]. In the first case, we have photonic crystals in space, which exhibit a band structure in the frequency domain. In the second case, we deal with photonic time crystals, where the wave number domain reveals a band structure. In the optical regime, the experimental realization of photonic time crystals is a formidable challenge because the required modulation depth of the refractive index must be significant, and the frequency of its changes comparable to optical frequencies [4].
@@
In this Letter, we pave the way for optical timetronics by demonstrating that a simple traveling wave resonator can exhibit a broad range of condensed matter phases or combinations of different phases, observed in the time domain. For example, phases like Anderson or topological insulators can be realized, different behaviors can be connected together, external fields can be realized and can act during a certain phase or the entire experiment, and the whole system can be completely reconfigured at any moment during the experiment. All of this is possible through time periodic modulation of the permittivity adapted to the shape of the small fragment of the resonator where the modulation is performed. That is, temporal harmonics of the modulation match resonantly with spatial harmonics of the resonator fragment and determine the effective behavior of the system. The described time modulation is experimentally feasible and allows for processing electromagnetic signals using phenomena known in condensed matter physics, thus paving the way for timetronics, i.e., applications where crystalline structures in time play a crucial role [38]. We present our idea by providing a detailed prescription for designing an arbitrary one-dimensional band structure. We then present two specific examples: the Su–Schrieffer–Heeger (SSH) model [39, 40] and the Wannier-Stark ladder [41, 42].
As an illustration of our idea, let us consider a closed resonator in the form of a ring with a square cross-section [Fig. 1(a)]. For simplicity, we re-scale the magnetic field vector, $\mathbf{H}\rightarrow\sqrt{\varepsilon_{0}/\mu_{0}}\,\mathbf{H}$, and use $L/2\pi$ and $L/2\pi c$ as the units of length and time, respectively, where $L$ is the circumference of the ring, $c$ is the speed of light, and $\mu_{0}$ and $\varepsilon_{0}$ are the vacuum permeability and permittivity [43]. Most importantly, we assume that the relative permittivity of a material $\varepsilon$ is constant everywhere in the resonator except for a small segment where $\varepsilon$ changes periodically in time according to
@@
for example, as in Fig. 1(b). Here $z$ denotes the position along the resonator and $h(z)$ and $f(t)$ are functions which vary in space and time respectively and describe the small segment of the resonator. In numerical calculations we set $\varepsilon_{r}=4$.
Figure 1: (a) Resonator in the form of a closed ring with a square cross-section. The circumference of the ring, in the units used in the text, is $2\pi$. In a small segment of the resonator described by $h(z)$, the permittivity is periodically modulated in time with the frequency $\Omega=2\pi/T$. (b) Example of the periodic modulation, $f(t)$, of the permittivity (left) in a small fragment of the resonator described by $h(z)$ of the Gaussian shape (right). This example is analyzed in the text. (c) Dispersion relation (in the laboratory frame) for longitudinal modes of the resonator with the length of the side of the cross-section area of the resonator $a=0.54$. Points indicate the discrete spectrum of the ring-shaped resonator. (d) In the frame moving with the frequency $\Omega$ which matches the free spectral range of the resonator, the dispersion relation for the longitudinal modes has a minimum at $k_{0}=40$, and superpositions of waves with $k\approx k_{0}$ evolve extremely slowly.
Next, we assume that the TE11 mode of the electromagnetic field has been injected into the resonator [43]. In this case, the electric field does not have a longitudinal component and we only need to consider the dependence of the transverse electric field amplitude $E(z,t)$ on time and space. Since $\varepsilon(z,t)$ is a periodic function of time, we can apply the Floquet theorem [44, 28] and seek solutions of Maxwell equations in the form $E(z,t)={\widetilde{E}}(z,t)e^{i\omega t}$ with ${\widetilde{E}}(z,t+T)={\widetilde{E}}(z,t)$. Here $\omega$ denotes the quasi-frequency of the electromagnetic field. The general solution of Maxwell equations can be obtained as a superposition of solutions ${\widetilde{E}}(z,t)e^{i\omega t}$. When solving Maxwell equations, it is convenient not to fully eliminate the magnetic field and to reduce them instead to a generalized eigenvalue problem
@@
with eigenvalue $\omega$ [43]. In this equation, $H(z,t)={\widetilde{H}}(z,t)e^{i\omega t}$, with ${\widetilde{H}}(z,t+T)={\widetilde{H}}(z,t)$, is the amplitude of the longitudinal component of the magnetic field and $k_{\perp}=\pi/a$ with $a$ being the normalized length of each side of the cross-sectional area of the resonator [Fig. 1(a)].
If there is no modulation in time of the permittivity (i.e., if $f(t)=0$), the solutions of Maxwell equations are characterized by a nonlinear dispersion relation $\omega=\sqrt{2k_{\perp}^{2}+k^{2}}$ where the wave number $k$ of the longitudinal modes takes integer values because we are considering the resonator in the form of a closed ring with periodic boundary conditions (Fig. 1). Let us assume that the permittivity is periodically modulated in time with the frequency $\Omega=2\pi/T$ and that the resonance condition with the frequency of a wave packet circulation around the resonator is satisfied, i.e., $\Omega$ matches the free spectral range of the resonator. In other words there exists a wave number $k_{0}$ for which the group velocity (calculated in the absence of the modulation) $\partial\omega/\partial k|_{k=k_{0}}=\Omega$ where the new units have been applied. For the sufficiently long resonator, the frequency $\Omega$ of the permittivity modulation is within the experimentally achievable range, e.g., for $L=1$ cm, $\Omega=15$ GHz.
To simplify the description of the resonant behavior of the system, we will switch to a reference frame evolving with the modulation frequency $z^{\prime}=z-\Omega t$ using the transformation $U=e^{\Omega t\partial_{z}}$ [43]. In the moving frame, the dispersion relation has a quadratic behavior with the minimum at the resonant wave number $k_{0}$, i.e. $\omega\approx\omega(k_{0})+\partial^{2}\omega/\partial k^{2}|_{k=k_{0}}(k-k_{0})^{2}/2$, see Fig. 1(d). In the present work we assume that the group velocity dispersion results from the geometry of the resonator but in general there can also be a contribution from the material properties of the resonator. In the moving frame, the evolution of a superposition of waves with wave numbers $k\approx k_{0}$ proceeds very slowly, and we can average Maxwell equations over time. This is essentially the rotating wave approximation, leading to time-independent effective Maxwell equations in the moving frame [45, 28, 43]:
@@
For clarity, we omitted the primes indicating that all variables and quantities refer to the moving frame. Moreover, the time-averaged permittivity $\bar{\varepsilon}(z)$ in the above equation can be written as
@@
In the moving frame, the effective Maxwell equations (3) describe electromagnetic waves in a resonator with a time-averaged permittivity $\bar{\varepsilon}(z)$ which varies in space along the resonator. In the laboratory frame, the segment of the resonator in which the permittivity is modulated in time is described by a localized function $h(z)$ and, therefore, has many non-zero Fourier coefficients $h_{m}$. The Fourier expansion coefficients of $f(t)$ can be chosen to realize any average permittivity $\bar{\varepsilon}(z)$ in the effective description in the moving frame since $\bar{\varepsilon}(z)$ can always be expanded such that $\bar{\varepsilon}(z)=\varepsilon_{r}+\sum_{m}\varepsilon_{m}e^{imz}$. All we need to do to obtain the given $\bar{\varepsilon}(z)$ in (4) is to choose $f(t)$ such that its Fourier coefficients satisfy $f_{-m}=\varepsilon_{m}/h_{m}$. For example, we can realize $\bar{\varepsilon}(z)\propto\cos(sz)$ with integer $s\gg 1$ and observe a band structure in the quasi-frequency, $\omega$, domain. We can realize a topological insulator or introduce disorder in a crystalline structure and realize a Anderson insulator. We can also combine crystalline structures with different properties, e.g., in different regions of $z$, the average permittivity $\bar{\varepsilon}(z)$ can reveal topologically different structures. Static electric field like potential, barriers or wells or more complex behavior can be realized in $\bar{\varepsilon}(z)$. Note that a stationary solution of (3) will appear as a propagating solution when we return to the laboratory frame. Thus, any condensed matter like behavior which we realize and observe versus $z$ in the moving frame will be observed in the time domain if, in the laboratory frame, we place a detector at a certain position $z_{0}$ in the resonator and investigate its clicking in time. This paves the way for optical timetronics, i.e., similar to electronics, we can design time-varying optical systems where electromagnetic signals are processed employing phenomena known in solid-state physics [38].
@@
Figure 2: (a) Quasi-frequency spectra corresponding to the exact and effective Maxwell equations. In the effective equations, the averaged permittivity is described by (5), where $s=12$ and $\lambda_{2}=2\pi\times 10^{-3}$. (b) The same as in (a), but in the presence of additional time modulation of the permittivity, which leads to a Gaussian barrier in (5) with a width of $\sqrt{2}\pi/41$ and $\lambda_{b}=-4.3\times 10^{-2}$ located at $z=0$. For $\lambda_{1}<0$, the formation of degenerate levels in the gap of the quasi-frequency bands can be observed. (c) Variations of the electric field at position $z_{0}=\pi$ in the laboratory frame corresponding to two quasi-frequency levels located in the gap between two bands in (b) for $\lambda_{1}/\lambda_{2}=-4$. The field is localized around the moment in time when a Gaussian barrier in $\bar{\varepsilon}(z_{0}-\Omega t)$ appears at the position $z_{0}$ [46]. (d) Similar to (c) but for a state from one of the bands. Such bulk states are delocalized along the entire period $T$. In all panels, both the exact results and the results of the effective Maxwell equations are presented.
@@
As a first example, we consider a system that can reveal topologically protected edge states in the time domain [47, 48, 28, 30, 49, 50]. Suppose $k_{\perp}=5.74$, $h(z)=e^{-z^{2}/2\sigma^{2}}$ with $\sigma=\pi/41$ and $f(t)=(\lambda_{1}/h_{s/2})\cos(s\Omega t/2)+(\lambda_{2}/h_{s})\cos(s\Omega t)$ with $s$ being even, with $\Omega=0.49$ and with $h_{s/2}$ and $h_{s}$ denoting the Fourier coefficients of $h(z)$, see Fig. 1(b). In the moving frame, the resulting average permittivity takes the form of a crystalline structure in space with a two-point basis,
@@
The quasi-frequency spectra obtained for $s=12$ with the help of the effective approach (3) and by solving the Maxwell equations (2) exactly are presented in Fig. 2(a). Both solutions match each other very well. For example, for $\lambda_{1}=0$, we observe a single band in the quasi-frequency domain consisting of $s=12$ levels. When $\lambda_{1}$ starts to differ from zero, the crystalline structure in $\bar{\varepsilon}(z)$ has a two-point basis and the initially single band splits into two. When we focus on these two bands, the band structure of the system is equivalent to the well-known SSH model which reveals topologically protected edge states in the presence of edges and in the topologically nontrivial regime [39, 40]. The crystalline structure in (5) is a periodic structure in a resonator without any edge. However, by a proper additional time modulation of the permittivity we can introduce a localized barrier in the crystalline structure and thus an edge in the system. Indeed, by introducing an additional modulation to our chosen $f(t)$ in the form of $\lambda_{b}\,e^{-t^{2}\Omega^{2}/2\sigma^{2}}$ we create a barrier in $\bar{\varepsilon}(z)$ represented by a Gaussian function with a width of $\sqrt{2}\pi/41$. In this case, we observe the appearance of two levels in the quasi-frequency spectrum located in the gap if $\lambda_{1}<0$ [Fig. 2(b)]. In the moving frame, the electromagnetic fields corresponding to these levels are localized on either side of the barrier, i.e. they are topologically protected edge states [40]. In the laboratory frame, if we place a detector at a certain position $z_{0}$ in the resonator, we can observe these edge states in the time domain [47, 48, 28, 30, 49]. This means, temporal changes of the probability of the detector clicking reveal the appearance of an edge state in time when the edge in $\bar{\varepsilon}(z_{0}-\Omega t)$ reaches the detector position. In Figs. 2(c)-2(d), we present how the electric field changes in time at the detector position for both the edge states and the so-called bulk states. The latter are delocalized in time across the entire period $T=2\pi/\Omega$. Note that the modulation depth of the permittivity needed to realize the described phenomena is very small, i.e. maximally of the order $10^{-2}$.
@@
Figure 3: (a) Quasi-frequency spectrum as a function of the strength, $\lambda$, of the artificial static electric field generated in (5) by the additional time modulation of the permittivity (see text) for $s=6$, $\lambda_{1}=0$ and $\lambda_{2}=4\times 10^{-3}$. (b) In the presence of the artificial static electric field, five solutions of the Maxwell equations exhibit Stark localization which in the laboratory frame we observe in the time domain. Namely, the electromagnetic fields at a fixed position in the laboratory frame (here $z_{0}=\pi$) are localized at different moments in time. These moments correspond to different local minima of $\bar{\varepsilon}(z_{0}-\Omega t)$. There is one more solution (green curve) localized around the discontinuity in $\bar{\varepsilon}(z_{0}-\Omega t)$ corresponding to the third quasi-frequency level in (a). The presented states correspond to $\lambda/\lambda_{2}=0.125$.
@@
As a second example, we consider the introduction of an artificial static electric field in the crystalline structure described by (5) and predict the observation of the Wannier-Stark localization in the optical system [41, 42]. Knowing the Fourier expansion of the linear potential in a resonator, $\lambda(z-\pi)=i\lambda\sum_{m\neq 0}e^{imz}/m$, we can introduce additional Fourier coefficients in $f(t)$, i.e. $f_{-m}=i\lambda/(mh_{m})$, which lead to a tilted crystalline structure and thus the presence of an artificial static electric field in our system. Note that in a ring-shaped resonator, the linear potential has a discontinuity at $z=0$ (or equivalently $z=2\pi$). However, away from $z=0$, the influence of the boundary conditions is negligible, and we can observe states that lose the character of extended Bloch waves and localize in different cells of the tilted crystalline structure $\bar{\varepsilon}(z)$. In the case of $s=6$, which is presented in Fig. 3, we see five such states with equally spaced quasi-frequencies. There is one more state localized around the discontinuity whose quasi-frequency falls between the second and third equally spaced eigenvalues [Fig. 3(a)]. In the laboratory frame, with a detector at a certain position $z_{0}$ in the resonator, we can observe the electromagnetic field localized at different moments in time [Fig. 3(b)].
In summary, we demonstrate that the properties of electromagnetic waves propagating in a resonator with a time-modulated refractive index can exhibit a range of behaviors known from research of crystalline structures in space, but observed in the time domain. Furthermore, different time crystalline structures can be combined by choosing the appropriate time modulation. Importantly, the realization of these structures in the optical regime does not require either deep modulation of the refractive index or modulation frequencies comparable to optical frequencies, as is the case in already known photonic time crystals. Thus, we obtain a tool that allows for processing of electromagnetic signals where time crystalline structures play a key role and can have practical applications. This paves the way for optical timetronics. As an example, in this Letter we considered a resonator in the shape of a one-dimensional ring due to the simplicity of their theoretical description. In our analysis, we have quantitatively examined the necessary conditions to realize the Su–Schrieffer–Heeger and Wannier-Stark ladder models, but our findings extend well beyond these particular cases. Notably, the introduction of a non-linear medium, where the permittivity varies with the electric field strength, enables us to introduce interactions. This capability opens up new opportunities to explore and study a wide range of condensed matter phases and phenomena.
@@
Supplemental Material: Towards Timetronics with Photonic Systems

### v2

@@ Abstract
Periodic driving of particles can create crystalline structures in their dynamics. Such systems can be used to study solid-state physics phenomena in the time domain. In addition, it is possible to realize photonic time crystals and to engineer the wave-number band structure of optical devices by periodic temporal modulation of the properties of light-propagating media. Here we introduce a versatile approach which uses traveling wave resonators to emulate various condensed matter phases in the time dimension. This is achieved by utilizing temporal modulation of the permittivity and the shape of small segments of the resonators. The required frequency and depth of the modulation are experimentally achievable which opens a pathway for the practical realisation of crystalline structures in time in microwave and in optical systems.
@@
Now, let us consider electromagnetic waves propagating in media whose refractive index changes periodically in space or in time [35, 36, 37]. In the first case, one can observe photonic crystals in space which exhibit a band structure in the frequency domain. In the second case, photonic time crystals emerge, and the wave number domain reveals a band structure. In the optical regime, the experimental realization of photonic time crystals is a formidable challenge because the required modulation depth of the refractive index must be significant. In addition, the frequency of refractive index changes must be comparable to optical frequencies [38].
In this Letter, we pave the way for optical timetronics by demonstrating that simple traveling wave resonators can exhibit a broad range of condensed matter phases, including combinations of different phases, in the time domain. For example, Anderson or topological insulators can be realized and different behaviors can be connected together. Moreover, experiments can be controlled by external fields which can be present during a certain phase or during the entire experiment, and which can be completely reconfigured at any moment. All of this is possible through time periodic modulation of the permittivity adapted to the shape of small fragments of the resonators where the modulation is performed. That is, temporal harmonics of the modulation match resonantly with spatial harmonics of the resonator fragments and determine the effective behavior of the system. The proposed time modulation is experimentally feasible and allows for processing electromagnetic signals using phenomena known in condensed matter physics, thus introducing a path towards timetronics, i.e., applications where crystalline structures in time play a crucial role [34].
In the following, we present our idea by providing a detailed prescription for designing an arbitrary one-dimensional band structure. We then present two specific examples: the Su–Schrieffer–Heeger (SSH) model [39, 40] and the Wannier-Stark ladder [41, 42]. In [43], we describe how to generalize our idea to two-dimensional condensed matter phases in periodically modulated coupled traveling-wave resonators.
As an illustration of our idea, let us consider a closed resonator in the form of a ring with a square cross-section [Fig. 1(a)]. For simplicity, we re-scale the magnetic field vector, $\mathbf{H}\rightarrow\sqrt{\varepsilon_{0}/\mu_{0}}\,\mathbf{H}$, and use $L/2\pi$ and $L/2\pi c$ as the units of length and time, respectively, where $L$ is the circumference of the ring, $c$ is the speed of light, and $\mu_{0}$ and $\varepsilon_{0}$ are the vacuum permeability and permittivity [43]. Most importantly, we assume that the relative permittivity of a material $\varepsilon$ is constant everywhere in the resonator except for a small segment where $\varepsilon$ changes periodically in time, for example such that
@@
as in Fig. 1(b). Here $z$ denotes the position along the resonator and $h(z)$ and $f(t)$ are functions which vary in space and in time respectively and describe the small segment of the resonator. In numerical calculations we set $\varepsilon_{r}=4$.
Figure 1: (a) Resonator in the form of a closed ring with a square cross-section. The circumference of the ring, in the units used in the text, is $2\pi$. In a small segment of the resonator described by $h(z)$, the permittivity is periodically modulated in time with the frequency $\Omega=2\pi/T$. (b) Example of the periodic modulation, $f(t)$, of the permittivity (left) in a small fragment of the resonator described by $h(z)$ of the Gaussian shape (right). This example is analyzed in the text. (c) Dispersion relation (in the laboratory frame) for longitudinal modes of the resonator with the length of the side of the cross-section area of the resonator $a=0.014\pi$. Points indicate the discrete spectrum of the ring-shaped resonator. (d) In the frame moving with the frequency $\Omega$ which matches the free spectral range of the resonator, the dispersion relation for the longitudinal modes has a minimum at $k_{0}=40$, and superpositions of waves with $k\approx k_{0}$ evolve extremely slowly.
Next, we assume that the TE11 mode of the electromagnetic field has been injected into the resonator [43]. In this case, the electric field does not have a longitudinal component and we only need to consider the dependence of the transverse electric field amplitude $E(z,t)$ on time and on space. Since $\varepsilon(z,t)$ is periodic in time, we can apply the Floquet theorem [44, 24] and seek solutions of Maxwell equations in the form $E(z,t)={\widetilde{E}}(z,t)e^{i\omega t}$ with ${\widetilde{E}}(z,t+T)={\widetilde{E}}(z,t)$. Here $\omega$ denotes the quasi-frequency of the electromagnetic field. The general solution of Maxwell equations can be obtained as a superposition of solutions ${\widetilde{E}}(z,t)e^{i\omega t}$. When solving Maxwell equations, it is convenient not to fully eliminate the magnetic field and to reduce them instead to the generalized eigenvalue problem
@@
with eigenvalue $\omega$ [43]. In this equation, $H(z,t)={\widetilde{H}}(z,t)e^{i\omega t}$, where ${\widetilde{H}}(z,t+T)={\widetilde{H}}(z,t)$ is the amplitude of the longitudinal component of the magnetic field and $k_{\perp}=\pi/a$ with $a$ being the length of each side of the cross-sectional area of the travelling wave resonator [Fig. 1(a)].
When the permittivity of the resonator is constant in time (i.e., when $f(t)=0$), the solutions of Maxwell equations are characterized by a nonlinear dispersion relation $\omega=\sqrt{2k_{\perp}^{2}+k^{2}}$ where the wave number $k$ of the longitudinal modes takes integer values because we are considering a closed ring resonator with periodic boundary conditions (Fig. 1). Next let us assume that the permittivity is periodically modulated in time with a non-zero frequency $\Omega=2\pi/T$ and that the resonance condition for the frequency of a wave packet circulating around the resonator is satisfied, i.e., $\Omega$ matches the free spectral range of the resonator. In this case, there exists a wave number $k_{0}$ for which the group velocity (calculated in the absence of the modulation) $\partial\omega/\partial k|_{k=k_{0}}=\Omega$. Here the new units have been applied. For a sufficiently long resonator, the frequency $\Omega$ of the permittivity modulation is within the experimentally achievable range, e.g., for $L=1$ cm, we have $\Omega=150$ MHz.
To simplify the description of the resonant behavior of the system, we will switch to a reference frame evolving with the modulation frequency $z^{\prime}=z-\Omega t$ by using the transformation $U=e^{\Omega t\partial_{z}}$ [43]. In the moving frame, the dispersion relation has a quadratic behavior with the minimum at the resonant wave number $k_{0}$, i.e. $\omega\approx\omega(k_{0})+\partial^{2}\omega/\partial k^{2}|_{k=k_{0}}(k-k_{0})^{2}/2$, see Fig. 1(d) and [43]. In the present work we assume that the group velocity dispersion results from the geometry of the resonator but in general there can also be a contribution from the material properties of the resonator. In the moving frame, a superposition of waves with wave numbers $k\approx k_{0}$ evolves very slowly, and we can average Maxwell equations over time. This is essentially the rotating wave approximation, leading to time-independent effective Maxwell equations in the moving frame [45, 24, 43]:
@@
For clarity, we omitted the primes; all variables and quantities refer to the moving frame. Moreover, the time-averaged permittivity $\bar{\varepsilon}(z)$ in the above equation can be written as
@@
In the moving frame, the effective Maxwell equations (3) describe electromagnetic waves in a resonator with a time-averaged permittivity $\bar{\varepsilon}(z)$ which varies in space along the resonator. In the laboratory frame, the segment of the resonator in which the permittivity is modulated in time is described by a localized function $h(z)$ and, therefore, has many non-zero Fourier coefficients $h_{m}$. The Fourier expansion coefficients of $f(t)$ can be chosen to realize any average permittivity $\bar{\varepsilon}(z)$ in the effective description in the moving frame since $\bar{\varepsilon}(z)$ can always be expanded such that $\bar{\varepsilon}(z)=\varepsilon_{r}+\sum_{m}\varepsilon_{m}e^{imz}$. All we need to do to obtain the given $\bar{\varepsilon}(z)$ in (4) is to choose $f(t)$ such that its Fourier coefficients satisfy $f_{-m}=\varepsilon_{m}/h_{m}$. For example, we can realize $\bar{\varepsilon}(z)\propto\cos(sz)$ with integer $s\gg 1$ and observe a band structure in the quasi-frequency, $\omega$, domain. We can realize a topological insulator or introduce disorder in a crystalline structure and realize an Anderson insulator.
We can also combine crystalline structures with different properties, e.g., in different regions of $z$, the average permittivity $\bar{\varepsilon}(z)$ can reveal topologically different structures. Static electric fields, such as potential barriers or wells, can be applied or a more complex modulation of $\bar{\varepsilon}(z)$ can be realized. Note that a stationary solution of (3) will appear as a propagating solution when we return to the laboratory frame. Thus, any condensed matter like behavior which we realize and observe versus $z$ in the moving frame will be observed in the time domain, if we place a detector in the laboratory frame at a certain position $z_{0}$ along the resonator and investigate its clicking in time. This approach paves the way for optical timetronics, i.e., similar to electronics, we can design time-varying systems where electromagnetic signals are processed by employing phenomena known in solid-state physics [34].
@@
Figure 2: (a) Quasi-frequency spectra corresponding to the exact and effective Maxwell equations. In the effective equations, the averaged permittivity is described by (5), where $s=12$ and $\lambda_{2}=10^{-6}$. (b) The same as in (a), but in the presence of additional time modulation of the permittivity, which leads to a Gaussian barrier in (5) with a width of $\pi/16$ located at $z=0$. For $\lambda_{1}<0$, the formation of degenerate levels in the gap of the quasi-frequency bands can be observed. (c) Variations of the electric field at position $z_{0}=0$ in the laboratory frame corresponding to two quasi-frequency levels located in the gap between two bands in (b) for $\lambda_{1}/\lambda_{2}=-1$. The field is localized around the moment in time when a Gaussian barrier in $\bar{\varepsilon}(z_{0}-\Omega t)$ appears at the position $z_{0}$ [46]. (d) Similar to (c) but for a state from one of the bands. Such bulk states are delocalized along the entire period $T$. In all panels, both the exact results and the results of the effective Maxwell equations are presented.
@@
As a first example, we consider a system that can reveal topologically protected edge states in the time domain [47, 48, 24, 26, 49, 50]. Suppose $k_{\perp}=72.15$, $h(z)=e^{-z^{2}/2\sigma^{2}}$ with $\sigma=\pi/41$ and $f(t)=(\lambda_{1}/h_{s/2})\cos(s\Omega t/2)+(\lambda_{2}/h_{s})\cos(s\Omega t)$ with $s$ being even, with $\Omega=4.9\times 10^{-3}$ and with $h_{s/2}$ and $h_{s}$ denoting the Fourier coefficients of $h(z)$, see Fig. 1(b). In the moving frame, the resulting average permittivity takes the form of a crystalline structure in space with a two-point basis,
@@
The quasi-frequency spectra obtained for $s=12$ with the help of the effective approach (3) and by solving Maxwell equations (2) exactly are presented in Fig. 2(a). In an experiment, the number of lattice sites can be increased either by increasing the modulation frequency $s\Omega$ or by extending the circumference of the ring resonator. Both the effective and exact solutions match each other very well. For example, for $\lambda_{1}=0$, we observe a single band in the quasi-frequency domain consisting of $s=12$ levels. When $\lambda_{1}$ starts to differ from zero, the crystalline structure in $\bar{\varepsilon}(z)$ has a two-point basis and the initially single band splits into two. When we focus on these two bands, the band structure of the system is equivalent to the well-known SSH model which reveals topologically protected edge states in the presence of edges and in the topologically nontrivial regime [39, 40].
The crystalline structure in (5) is a periodic structure in a resonator without any edge. However, by a proper additional time modulation of the permittivity we can introduce a localized barrier in the crystalline structure and thus an edge in the system. Indeed, by introducing an additional modulation to our chosen $f(t)$ in the form of $\lambda_{b}\,e^{-t^{2}\Omega^{2}/2\sigma^{2}}$, where $\lambda_{b}=-9.8\times 10^{-6}$, we create a barrier in $\bar{\varepsilon}(z)$ represented by a Gaussian function with a width of $\pi/16$. In this case, we observe the appearance of two levels in the quasi-frequency spectrum located in the gap if $\lambda_{1}<0$ [Fig. 2(b)]. In the moving frame, the electromagnetic fields corresponding to these levels are localized on either side of the barrier, i.e. they are topologically protected edge states [40]. In the laboratory frame, if we place a detector at a certain position $z_{0}$ in the resonator, we can observe these edge states in the time domain [47, 48, 24, 26, 49]. This means, temporal changes of the probability of the detector clicking reveal the appearance of an edge state in time when the edge in $\bar{\varepsilon}(z_{0}-\Omega t)$ reaches the detector position. Figures 2(c)-2(d) show how the electric field changes in time at the detector position for both the edge states and the so-called bulk states. The latter are delocalized in time across the entire period $T=2\pi/\Omega$. In this experiment, a pulse with the central wavenumber $k_{0}$ and temporal width of $0.119/\Omega$ is injected at a specific position $z$. If the injection occurs precisely when the edge in $\bar{\varepsilon}(z-\Omega t)$ passes by, the pulse will propagate without any distortion, due to the topological properties of the system. In contrast, if the pulse is injected at other times, it will quickly spread over the entire resonator within a characteristic timescale of the order $s/\Omega$. Note that the depth of the permittivity modulation needed to realize the described phenomena is very small, i.e. maximally of the order $10^{-5}$.
@@
Figure 3: (a) Quasi-frequency spectrum as a function of the strength, $\lambda$, of the artificial static electric field generated in (5) by the additional time modulation of the permittivity (see text) for $s=6$, $\lambda_{1}=0$ and $\lambda_{2}=10^{-6}$. (b) In the presence of the artificial static electric field, five solutions of the Maxwell equations exhibit Stark localization which in the laboratory frame we observe in the time domain. Namely, the electromagnetic fields at a fixed position in the laboratory frame (here $z_{0}=0$) are localized at different moments in time. These moments correspond to different local minima of $\bar{\varepsilon}(z_{0}-\Omega t)$. There is one more solution (green curve) localized around the discontinuity in $\bar{\varepsilon}(z_{0}-\Omega t)$ corresponding to the third quasi-frequency level in (a). The presented states correspond to $\lambda/\lambda_{2}=0.5$.
@@
As a second example, we consider the introduction of an artificial static electric field in the crystalline structure described by (5) and predict the observation of the Wannier-Stark localization in the optical system [41, 42]. Knowing the Fourier expansion of the linear potential in a resonator, $\lambda(z-\pi)=i\lambda\sum_{m\neq 0}e^{imz}/m$, we can introduce additional Fourier coefficients in $f(t)$, i.e. $f_{-m}=i\lambda/(mh_{m})$, which lead to a tilted crystalline structure and thus the presence of an artificial static electric field in our system. Note that in a ring-shaped resonator, the linear potential has a discontinuity at $z=0$ (or equivalently at $z=2\pi$). However, away from $z=0$, the influence of the boundary conditions is negligible, and we can observe states that lose the character of extended Bloch waves and localize in different cells of the tilted crystalline structure $\bar{\varepsilon}(z)$. In the case of $s=6$, which is presented in Fig. 3, we see five such states with equally spaced quasi-frequencies. There is one more state localized around the discontinuity whose quasi-frequency falls between the second and third equally spaced eigenvalues [Fig. 3(a)]. In the laboratory frame, with a detector at a certain position $z_{0}$ in the resonator, we can observe the electromagnetic field localized at different moments in time if electromagnetic pulses are injected into the resonator at proper moments [Fig. 3(b)].
Different time crystalline structures can be combined by choosing the appropriate time modulation. Importantly, the realization of these structures in the optical regime does not require either deep modulation of the refractive index or modulation frequencies comparable to optical frequencies, as is the case in already known photonic time crystals. Thus, we obtain a tool that allows for processing of electromagnetic signals where time crystalline structures play a key role. This approach has practical applications and introduces a path towards optical timetronics.
As examples, we considered a resonator in the shape of a one-dimensional ring in this Letter. We quantitatively examined the necessary conditions to realize the Su–Schrieffer–Heeger and Wannier-Stark ladder models, but our findings extend well beyond these particular cases. Our idea can be readily extended to two-dimensional, and possibly higher-dimensional, condensed matter phases implemented in periodically modulated coupled traveling-wave resonators. [43]. Notably, the introduction of a non-linear medium, where the permittivity varies with the electric field strength, enables us to introduce interactions [43]. If the energy (in frequency units) associated with the nonlinear terms is comparable to or smaller than the bandwidths of the quasi-frequencies shown, for instance, in Figs. 2-3, then nonlinearity will not destroy the crystalline structures but instead will provide the possibility to obtain qualitatively different solutions within those structures. This capability opens up new opportunities to explore and study a wide range of condensed matter phases and phenomena. Finally, in this work, we focus on time-periodic modulations that preserve time-reversal symmetry. Our formalism, however, can be extended to model systems with time reversal symmetry-breaking modulations, resulting in an asymmetric effective permittivity $\bar{\varepsilon}(z)$. This is analogous to quantum systems, where the presence of magnetic field breaks time-reversal symmetry. Such extensions open new opportunities for realizing photonic devices and synthetic gauge fields.
@@
I Supplementary Material
II Rotating Frame